Line AB has a slope of -6 and an x-intercept of 6. The y-value of point A is twice its x-value. Find the y-value of point A.
Answer by ankor@dixie-net.com(22740)

Line AB has a slope of -6 and an x-intercept of 6. The y-value of point A is twice its x-value. Find the y-value of point A.
:
We can use the slope equation (y2 - y1)/(x2 - x1) = m
:
Let the x intercept be x1, y1: (6, 0)
let x2 = x; y2 = 2x
We know m = -6
Find x:
;
(2x - 0)/(x - 6) = -6
:
2x = -6(x - 6)
:
2x = -6x + 36
:
2x + 6x = 36
:
8x = 36
:
x = 36/8
:
x = 4.5
:
y = 2x therefore y = 9 at point A